OpenStudy (anonymous):

WILL MEDAL & FAN. Solve for r, I have the answer but im confused on how it was gotten P= A/1+rt do I multiply by the denominator? Answer is: r= A-P/pt Posting this again cause I cant get anyone to help and this bump thing doesn’t seem to be working…

OpenStudy (anonymous):

\[P= \frac{ A }{ 1+rt}\]

OpenStudy (anonymous):

Actually, you can multiply by the denominator, but since it is 1, it can be ignored as any number divided by 1 equals itself by the Identity Property of Division. Also your answer of r=A-P/pt isn't exactly correct, but I think you made a typo and meant \[r = \frac{ A - p }{ t }\] since you seemed to have everything mixed up.

OpenStudy (anonymous):

ok so you multiplied by the t only?

OpenStudy (anonymous):

No, I subtracted both sides by A and divided both sides by t to solve for r.

OpenStudy (anonymous):

Sorry, Im a little slow with the all variable equations >.< Ok so the 1 is basically ignored because of the property so by subtracting the A you break it from being a fraction?

OpenStudy (anonymous):

Oh wait, I realized your typo in the problem that equals P. Hold on a second.

OpenStudy (anonymous):

Says in the answer key \[\frac{ A-P }{ Pt }\]

OpenStudy (anonymous):

2 uppercase P?? o__o

OpenStudy (anonymous):

Honestly, I also don't see how they found that.

OpenStudy (anonymous):

I did this: \[P = \frac{ A }{ 1 + rt }\] This is how to solve for R. Rewrite P and p/1 and use a proportion to get \[P(1 + rt) = A\] Then divide both sides by P to get \[1 + rt = \frac{ A }{ P }\] Subtract 1 from both sides then divide both sides by t. \[r = \frac{ \frac{ A }{ P } - 1 }{ t }\] Yes, it is strange but this is how I did it with my methods.

OpenStudy (anonymous):

yes it makes sense. so im guessing insead of the -1 they substitute it by another P

OpenStudy (anonymous):

Wait, I used a website and found this method:

OpenStudy (anonymous):

There is another answer choice of \[r=\frac{ P-A }{ 1+t }\]

OpenStudy (anonymous):

but it still says the right one is the one i said

OpenStudy (anonymous):

My mistake since I don't really always see other methods. But now I see the pathway you chose after seeing this on the website. \[P = \frac{ A }{ (1 + rt) }\] \[P = \frac{ A }{ rt + 1 }\] \[(rt + 1)P = A\] \[r = \frac{ a - P }{ Pt }\] Yes you are correct. Sorry I didn't see how you arranged the variables in that manner earlier.

OpenStudy (anonymous):

Ok so you divided by the P in the (rt+1)P=A

OpenStudy (anonymous):

But how is it that the r is left alone when its stuck to the t and +1

OpenStudy (anonymous):

@MathematicsNerd

OpenStudy (anonymous):

\[rtP + P = A\] Apply distributive property \[rtP = A - P\] Subtraction Property pf Equality \[r = \frac{ A - P }{ Pt }\] Divide both sides by Pt or tP